# Masked Diffusion for Generative Recommendation

### v1

@@ 4.1 Experimental Setup
To understand the performance of MADRec as we scale number of SIDs per item, we perform an ablation on changing number of SIDs while fixing the rest of the setting. We report our results in Table 3. We observe that the performance of MADRec improves as we increase number of SIDs from 3 to 4 but decreases as we go from 4 to 5. A plausible reason for the decrease in performance from 4 to 5 is the potential increase in invalid predicted SIDs, as the model may predict SIDs that do not correspond to a valid item. Combining MADRec with constrained beam search to prevent such invalid SIDs as we scale the number of SIDs would be an interesting future direction [hou2025generating].
@@
As MADRec training does not incorporate an inductive bias toward any particular token order – such as the left-to-right ordering used in autoregressive training – it offers greater flexibility in choosing which tokens to unmask during inference [zheng2024reparameterized, kim2025train]. In this section, we evaluate three strategies for selecting the tokens to unmask:
@@
Table 4: Performance of different inference strategies for MADRec. By default, MADRec uses the Greedy strategy.
@@
MADRec + Random 5.01 7.54 3.27 4.09
MADRec + Left-to-right 5.31 8.09 3.46 4.37
MADRec (Greedy) 5.38 8.15 3.51 4.41
@@
In this section, we present the experimental results of extending MADRec with dense retrieval. The MADRec with dense retrieval unifies the MADRec’s SID generation capabilities with dense retrieval as described in Section 3.3. We use the same 4096-dimensional Flan-T5-XXL embeddings that we used for SID assignment as the items’ text embeddings. We construct the predicted dense embedding of an item by combining the output of the 4th layer of our encoder model and projecting it to 4096 dimensions (matching the text embedding size) through a one-layer MLP with a hidden dimension of 256. Since both the text and predicted embeddings are high-dimensional, MLP with full-rank weights would introduce over one million parameters per layer. Therefore, we use an MLP with low-rank weights of rank $32$. During training, with a probability of $\beta=0.2$, we mask all SIDs of an item to promote learning at the item-level abstraction. In the unified retrieval setup—where SID generation and dense retrieval are jointly leveraged—MADRec first generates 20 beams using beam search and then re-ranks them based on their dense retrieval scores to obtain the top-10 candidates.
We report our results on the Beauty dataset in Table 5. The results indicate that integrating MADRec with dense retrieval and unified retrieval both improve performance compared to the generative retrieval baseline, while the two enhanced variants perform comparably to each other. Perhaps more importantly, these observations exemplify that MADRec is general enough to be compatible with auxiliary methods developed for improving performance in AR modeling for GR.
Table 5: Performance comparison of MADRec combined with different retrieval methods. By default, MADRec uses the generative retrieval strategy outlined in Section 3.2.
@@
MADRec 5.38 8.15 3.51 4.41
MADRec + Dense Retrieval 5.41 8.50 3.53 4.45
MADRec + Unified Retrieval 5.43 8.59 3.54 4.47
@@ 5 Related work
Since its introduction in VQ-REC [hou2023learning] and TIGER [rajput2023recommender], much of the work in GR with SIDs has focused on improving item tokenization, via methods such as contrastive learning [zhu2024cost], incorporating collaborative signals [xiao2025progressive, hua2023index, qu2025tokenrec, wang2024learnable, wang2024eager, liu2024end], directly using LLMs [tan2024idgenrec, jin2023language] and encouraging more uniform SID distributions [kuai2024breaking]. Other works have conditioned SID generation on user profiles [paischer2024preference], merged GR with dense retrieval [yang2024unifying, yang2025sparse], and employed tree-based decoding [feng2022recommender, si2024generative]. However, all of these works adopt the AR training strategy from [rajput2023recommender]. RPG [hou2025generating] proposes to use product-quantized SIDs to enable decoding the SID tokens for a single item in parallel, however this approach requires decoding items sequentially and needs many SIDs per item to see performance gains.
In contrast, our method allows for parallel decoding with any SIDs, achieving SOTA even with the standard residual-quantized SIDs. Outside of GR, [ren2024non] and [valluri2024scaling] employ non-AR, non-diffusion generative models for reranking and information retrieval, respectively.
@@ 6 Conclusion
We present a novel modeling paradigm for GR with SIDs inspired by recent advancements in masked diffusion modeling in NLP. Empirically, our framework brings several advantages over standard AR modeling, including improved generalization in data-sparse settings, inference efficiency, and coarse-grained recall. Our framework is also simple and general enough to incorporate innovations from AR SID modeling, such as incorporating dense retrieval. Exploring novel design choices to combine other such auxiliary techniques with MADRec presents an interesting direction for future research. We also envision that user sequence modeling with masked diffusion can be further improved by a more sophisticated training and inference guidance strategy (e.g., classifier-free/classifier-based guidance [schiff2024simple] or error correction via remasking [wang2025remasking, von2025generalized]).
@@ A.1 Additional Experimental Details
Implementation details. Our implementation of MADRec is based on the GRID codebase presented in [grid]. We use the AdamW optimizer with a learning rate of 0.005, a weight decay of 0.001, a batch size of 8192, and early stop all experiments using validation recall@10. We use the implementation of TIGER [rajput2023recommender] provided in GRID. We use the implementation of LIGER [yang2024unifying] using the public repository released by the paper’s authors. Both TIGER and LIGER use 4 SIDs per item, including the deduplication token, consistent with their original implementations and TIGER’s optimal setting for the Amazon datasets [grid]. All SIDs are assigned by executing residual $k$-means clustering with $c=256$ clusters per layer on mean-pooled Flan-T5-XXL embeddings. For baselines, we use SASRec results reported by Rajput et al. [rajput2023recommender], and BERT4Rec, DreamRec, and CaDiRec results from Cui et al. [cui2024cadirec].
@@ A.2 Additional Experimental Results
Table 7: Comparison of the performance of MADRec with other GR methods on multiple datasets.

### v2

@@ 4.1 Experimental Setup
To understand the performance of MaskGR as we scale number of SIDs per item, we perform an ablation on changing number of SIDs while fixing the rest of the setting. We report our results in Table 3. We observe that the performance of MaskGR improves as we increase number of SIDs from 3 to 4 but decreases as we go from 4 to 5. A plausible reason for the decrease in performance from 4 to 5 is the potential increase in invalid predicted SIDs, as the model may predict SIDs that do not correspond to a valid item. Combining MaskGR with constrained beam search to prevent such invalid SIDs as we scale the number of SIDs would be an interesting future direction [hou2025generating].
@@
As MaskGR training does not incorporate an inductive bias toward any particular token order – such as the left-to-right ordering used in autoregressive training – it offers greater flexibility in choosing which tokens to unmask during inference [zheng2024reparameterized, kim2025train]. In this section, we evaluate three strategies for selecting the tokens to unmask:
@@
Table 4: Performance of different inference strategies for MaskGR. By default, MaskGR uses the Greedy strategy.
@@
MaskGR + Random 5.01 7.54 3.27 4.09
MaskGR + Left-to-right 5.31 8.09 3.46 4.37
MaskGR (Greedy) 5.38 8.15 3.51 4.41
@@
In this section, we present the experimental results of extending MaskGR with dense retrieval. The MaskGR with dense retrieval unifies the MaskGR’s SID generation capabilities with dense retrieval as described in Section 3.3. We use the same 4096-dimensional Flan-T5-XXL embeddings that we used for SID assignment as the items’ text embeddings. We construct the predicted dense embedding of an item by combining the output of the 4th layer of our encoder model and projecting it to 4096 dimensions (matching the text embedding size) through a one-layer MLP with a hidden dimension of 256. Since both the text and predicted embeddings are high-dimensional, MLP with full-rank weights would introduce over one million parameters per layer. Therefore, we use an MLP with low-rank weights of rank $32$. During training, with a probability of $\beta=0.2$, we mask all SIDs of an item to promote learning at the item-level abstraction. In the unified retrieval setup—where SID generation and dense retrieval are jointly leveraged—MaskGR first generates 20 beams using beam search and then re-ranks them based on their dense retrieval scores to obtain the top-10 candidates.
We report our results on the Beauty dataset in Table 5. The results indicate that integrating MaskGR with dense retrieval and unified retrieval both improve performance compared to the generative retrieval baseline, while the two enhanced variants perform comparably to each other. Perhaps more importantly, these observations exemplify that MaskGR is general enough to be compatible with auxiliary methods developed for improving performance in AR modeling for GR.
Table 5: Performance comparison of MaskGR combined with different retrieval methods. By default, MaskGR uses the generative retrieval strategy outlined in Section 3.2.
@@
MaskGR 5.38 8.15 3.51 4.41
MaskGR + Dense Retrieval 5.41 8.50 3.53 4.45
MaskGR + Unified Retrieval 5.43 8.59 3.54 4.47
@@ 5 Related work
Since its introduction in VQ-REC [hou2023learning] and TIGER [rajput2023recommender], much of the work in GR with SIDs has focused on improving item tokenization, via methods such as contrastive learning [zhu2024cost], incorporating collaborative signals [xiao2025progressive, hua2023index, qu2025tokenrec, wang2024learnable, wang2024eager, liu2024end], directly using LLMs [tan2024idgenrec, jin2023language] and encouraging more uniform SID distributions [kuai2024breaking]. Other works have conditioned SID generation on user profiles [paischer2024preference], merged GR with dense retrieval [yang2024unifying, yang2025sparse], and employed tree-based decoding [feng2022recommender, si2024generative]. However, all of these works adopt the AR training strategy from [rajput2023recommender]. RPG [hou2025generating] proposes to use product-quantized SIDs to enable decoding the SID tokens for a single item in parallel, however this approach requires decoding items sequentially and needs many SIDs per item to see performance gains. In contrast, our method allows for parallel decoding with any SIDs, achieving SOTA even with the standard residual-quantized SIDs. Outside of GR, [ren2024non] and [valluri2024scaling] employ non-AR, non-diffusion generative models for reranking and information retrieval, respectively.
@@ 6 Conclusion
We present a novel modeling paradigm for GR with SIDs inspired by recent advancements in masked diffusion modeling in NLP. Empirically, our framework brings several advantages over standard AR modeling, including improved generalization in data-sparse settings, inference efficiency, and coarse-grained recall. Our framework is also simple and general enough to incorporate innovations from AR SID modeling, such as incorporating dense retrieval. Exploring novel design choices to combine other such auxiliary techniques with MaskGR presents an interesting direction for future research. We also envision that user sequence modeling with masked diffusion can be further improved by a more sophisticated training and inference guidance strategy (e.g., classifier-free/classifier-based guidance [schiff2024simple] or error correction via remasking [wang2025remasking, von2025generalized]).
@@ A.1 Additional Experimental Details
Implementation details. Our implementation of MaskGR is based on the GRID codebase presented in [grid]. We use the AdamW optimizer with a learning rate of 0.005, a weight decay of 0.001, a batch size of 8192, and early stop all experiments using validation recall@10. We use the implementation of TIGER [rajput2023recommender] provided in GRID. We use the implementation of LIGER [yang2024unifying] using the public repository released by the paper’s authors. Both TIGER and LIGER use 4 SIDs per item, including the deduplication token, consistent with their original implementations and TIGER’s optimal setting for the Amazon datasets [grid]. All SIDs are assigned by executing residual $k$-means clustering with $c=256$ clusters per layer on mean-pooled Flan-T5-XXL embeddings. For baselines, we use SASRec results reported by Rajput et al. [rajput2023recommender], and BERT4Rec, DreamRec, and CaDiRec results from Cui et al. [cui2024cadirec].
@@ A.2 Additional Experimental Results
Table 7: Comparison of the performance of MaskGR with other GR methods on multiple datasets.